What is the mass in grams of 11.5 l of N2 at 48.2 degree Celsius and 582mmHg?
How many moles of NO2 gas have a volume of 1,311.4 at 8.1 degree Celsius and 2.8 atm pressure?
1
Expert's answer
2020-11-18T13:43:43-0500

1.PV= nRT.

582mmHg to atm = 582/760 = 0.76579atm T = 48.2+273 = 321.2K

n= PV/RT = 0.76579 × 11.5 / (0.0821 × 321.2) = 0.33mol

Mole = mass/molar mass

Mass = mole × molar mass = 0.33 × (14×2)

= 9.24g


2. n= PV/RT. T = 8.1 + 273 = 281.1 K

n= 2.8 × 1311.4 / (0.0821 × 281.1)

n= 159.11 mole